# Can We Treat Noisy Labels as Accurate?

### v1

@@ 4 EchoAlign
The EchoAlign framework tackles the challenge of noisy labels in supervised learning. It comprises two primary components: (1) EchoMod modifies data instances to align them better with their potentially incorrect labels. (2) EchoSelect selects the most reliable modified instances, creating a cleaner, more consistent training dataset, which is used to train a general neural network.
@@ 4.1 EchoMod: Instance Modification
Motivation  When labels are noisy, they do not reflect the true characteristics of the corresponding data instances. This discrepancy can hinder a model’s ability to learn meaningful patterns. EchoMod addresses this by transforming data instances to be more consistent with their noisy labels. This controlled modification helps the model extract relevant information when the labels are noisy.
@@
Mechanism  EchoMod leverages a pre-trained controllable generative model (e.g., a controllable diffusion-based model) to modify data instances. The primary goal is to enhance the alignment between instances and their potentially noisy labels. This alignment is achieved by carefully guiding the generative model’s process. Firstly, the controllable generative model has undergone prior training on a large dataset. This pre-training has equipped the model with a deep understanding of the patterns and structures inherent in the data domain. Secondly, EchoMod provides both the original instance ($X$) and the noisy label ($\tilde{Y}$) as inputs to the generative model. This dual conditioning shapes the output, encouraging the model to produce a modified instance ($X^{\prime}$) that closely aligns with the noisy label while still retaining essential characteristics of the original data. Striking this balance between label alignment and preventing excessive distortion is crucial.
@@
Reasonableness, Effectiveness, and Fairness  Handling label noise in noisy label learning is a well-recognized challenge. Previous works have primarily focused on the utilization and optimization of internal data. Our approach introduces a novel perspective by integrating external knowledge to enhance model robustness. This integration does not compromise fairness, as our flexible framework accommodates various generative models, and can be fine-tuned for specific noisy label problems. By doing so, we ensure that the method does not overly rely on any particular model. This approach is particularly advantageous when dealing with ambiguous data. The inherent ambiguity in the data can lead to low confidence in direct discrimination. Instead, by modifying the input data through controllable generative models, we can better resolve discrepancies between instances and labels while preserving the meaningful characteristics of the original data. This not only improves the model’s discriminative capability but also enhances overall performance and reliability.
@@
No Fine-Tuning for Generalizability  In most cases, EchoMod can leverage a pre-trained controllable generative model without fine-tuning during the alignment process. This preserves the model’s ability to understand general data characteristics, promoting EchoAlign’s applicability across various domains. While EchoMod can be effective without fine-tuning, additional performance gains might be realized by tailoring the controllable generative model to highly specialized tasks or data distributions. In such cases, fine-tuning the parameters ($\theta$) could lead to better alignment between instances and noisy labels. Examples of this specialization include medical imaging or scientific data.
@@ 4.2 EchoSelect: Instance Selection
Motivation  While EchoMod improves the alignment between instances and noisy labels, some modified instances may still exhibit inconsistencies. Additionally, the instance modification process can introduce distribution shifts between the modified training data and the true test distribution. EchoSelect safeguards against these issues by identifying and retaining only the most reliable modified instances. This filtering enhances model robustness, reduces the impact of noisy data, and helps mitigate the distribution shifts caused by instance modification.
Mechanism  EchoSelect employs a metric to assess the similarity between modified instances and a reference representation of clean data. We use the Cosine similarity between feature vectors extracted using a suitable feature extractor (e.g., the image encoder of CLIP (Radford et al., 2021)):
@@
Selection Process  EchoSelect calculates similarity for all modified instances, comparing them to their original counterparts. To mitigate distribution shifts, priority is given to maintaining clean original instances as much as possible. The final training set consists of two parts: (1) Original instances with similarity larger than a determined threshold are deemed sufficiently aligned with clean data characteristics and retained, and (2) modified instances with similarity below the threshold are included. These instances are likely those where the modification was most beneficial in aligning them with the noisy labels, while also indicating some degree of difference from the original distribution. This threshold $\tau$ balances the inclusion of modified instances with the preservation of the original data distribution. This threshold ensures that only instances aligned with the characteristics of clean data are retained. Our sensitivity analysis (§5.3) verified the robustness of $\tau$ across types of noise.
@@ 4.3 EchoAlign: Optimized Combination
0:  Pre-trained controllable generative model $f_{\theta}$, Noisy dataset $(X,\tilde{Y})$, Threshold $\tau$, Feature Extractor
@@
1:  Generate modified instances: $X^{\prime}\leftarrow f_{\theta}(X,\tilde{Y})$
2:  Compute similarity: $S(X^{\prime},X)$ using Equation (1)
@@
The integration of EchoMod and EchoSelect enables the creation of a refined training dataset that is aligned with noisy labels and filtered for quality. This optimized dataset is better suited for robust learning in the presence of large label noise. Since the refined training dataset can be further used to train a supervised or self-supervised model for LNL, EchoAlign can be combined with advanced LNL methods to address the impact of label noises further. The integration of EchoMod and EchoSelect is encapsulated in the following Algorithm 1, which details the steps for modifying instances and selecting the optimal subset.
@@ 5.1 Experiment Setup
Dataset  Our experiments are conducted on two synthetic datasets: CIFAR-10, and CIFAR-100 (Krizhevsky et al., 2009), and a real-world dataset: CIFAR-10N (Wei et al., 2022). CIFAR-10 and CIFAR-100 each contain 50,000 training and 10,000 testing images, with a size of 32$\times$32, covering 10 and 100 classes respectively. CIFAR-10N utilizes the same training images from CIFAR-10 but with labels re-annotated by humans. Following previous research protocols (Bai et al., 2021; Xia et al., 2019, 2023b), we corrupted these synthetic datasets by three types of label noise. Specifically, symmetric noise randomly alters a certain proportion of labels to different classes to simulate random labeling errors; pair flip noise changes labels to specific adjacent classes with a certain probability; instance-dependent noise modifies labels based on the image features to related incorrect classes. Due to resource constraints, we use a light and representative human-annotated noisy dataset CIFAR-10N instead of Clothing1M (Xiao et al., 2015). A detailed runtime analysis is discussed in the Appendix D, demonstrating that the runtime is reasonable for all datasets used. For CIFAR-10N, we provide four noisy label sets: ‘Random i=1, 2, 3’, each representing the label provided by one of three independent annotators; and ‘Worst’, which selects the noisiest label when incorrect annotations are present.
@@
Baseline  We compare EchoAlign against various paradigms of baselines for addressing label noise. Under the robust loss function paradigm, we include APL (Ma et al., 2020), PCE (Menon et al., 2019), AUL (Zhou et al., 2023), and CELC (Wei et al., 2023) as baselines; under the loss correction paradigm, we utilize T-Revision (Xia et al., 2019) and Identifiability (Liu et al., 2023); under the label correction paradigm, we select Joint (Tanaka et al., 2018); and under the sample selection paradigm, we employ Co-teaching (Han et al., 2018), SIGUA (Han et al., 2020) and Co-Dis (Xia et al., 2023a). We compare these methods against a simple cross-entropy (CE) loss baseline. Following the fair baseline design used by Xia et al. (2023b), we do not compare with some methods such as MixUp (Zhang et al., 2018), DivideMix (Li et al., 2020), and M-correction (Arazo et al., 2019), as these involve semi-supervised learning, making such comparisons unfair due to inconsistent settings.
@@
Implementation Details  All experiments were conducted on an NVIDIA V100 GPU using PyTorch. The model architectures and parameter settings were kept consistent with previous studies (Bai et al., 2021). The experiments were configured with a learning rate of 0.1, using the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9, and weight decay set to $1\times 10^{-4}$. We applied 30% and 50% symmetric noise as well as 45% pair flip noise on the CIFAR-10 and CIFAR-100 datasets to assess model performance. The CIFAR-10 dataset utilized the standard ResNet-18 (He et al., 2016) architecture, while CIFAR-100 used ResNet-34. Similarly, for the CIFAR-10N dataset, the same ResNet-18 model was used. Prior to the training phase, ControlNet was utilized as our reference model in the controllable generation model module. This choice was strategic; ControlNet was the least effective model identified in prior analyses (Chen et al., 2023c). Employing this model underscores the robustness of our approach, ensuring that the efficacy of our method is not overly contingent upon the capabilities of any specific generative model. This decision highlights our method’s adaptability and general efficacy across varying scenarios. We employed the Canny edge detector as a simple preprocessor to extract features from the instances, using labels as textual controls with the prompt "a photo of {label}". This was executed without any additional or negative prompts, and the sampling process was limited to 20 steps. All experiments were repeated three times under different random seeds, and results are reported as averages with standard deviations.
5.2 Main Results
@@
We validated the effectiveness of our method on the CIFAR-10 and CIFAR-100 synthetic datasets. For each dataset, we utilized 90% of the samples with noisy labels for training and reserved 10% as validation samples, with testing conducted on clean samples. Part of the baseline results was taken from previous work (Xia et al., 2023b). As shown in Table 1, our results exhibited state-of-the-art performance across nearly all conditions on both datasets. Notably, under more challenging noise conditions, 50% instance-dependent noise on CIFAR-10 and 45% symmetric noise on CIFAR-100, our method significantly outperformed the current state-of-the-art baselines, which is mainly attributed to the robustness of our method to the noise rate. Importantly, our method demonstrates enhanced performance under higher noise rates, due to EchoMod’s noise-independence. This attribute allows the model to learn consistent instance features across different types of noise and noise rates. The primary cause of performance variation across different settings is the inherent difference in the number of clean samples within the datasets. Our approach also exhibited remarkable performance on the real-world dataset, CIFAR-10N, outperforming the state-of-the-art existing methods in all real noise conditions. Furthermore, our proposed method demonstrated exceptional robustness, with standard deviations in performance across all conditions within floating-point precision.
@@ 5.3 In-Depth Analyses
Ablation Analysis  To assess the effectiveness of Instance Modification and EchoSelect, we conducted ablation studies by systematically disabling these components. Specifically, we evaluated two configurations: Instance Modification Only and EchoSelect Only, and compared both against the standard Cross-Entropy Loss (CE) as a baseline. These experiments were carried out under several settings with high noise rates, which presented significant challenges for the model. The experimental results in Table 3 revealed that in the configuration using only Instance Modification, the model’s accuracy not only failed to exceed the baseline CE but also showed poorer performance. This decline primarily stems from the data distribution shift caused by solely using modified instances, which adversely affects the model’s generalization capability and performance. In contrast, while using only the EchoSelect component did improve performance, it still fell short compared to EchoAlign. This indicates that although EchoSelect significantly reduces the impact of noise, its effectiveness is limited by the available number of samples.

### v2

@@ 4 EchoAlign
The EchoAlign framework tackles the challenge of noisy labels in supervised learning. It comprises two primary components: (1) EchoMod modifies instances using controllable generative models, ensuring alignment with noisy labels while preserving intrinsic characteristics. (2) EchoSelect selects original instances with correct labels, maintaining a balanced distribution between original and modified data.
Figure 7: The framework of EchoAlign.
Figure 7 illustrates the overall framework of EchoAlign. Specifically, EchoMod utilizes controllable generative models (CGMs) to perform instance modifications based on noisy labels, while EchoSelect applies feature similarity evaluation to strategically filter instances, preserving original instances that are likely clean and adopting modified instances to ensure label alignment. The integration of these modules effectively addresses the characteristic and distribution shifts introduced by label noise.
@@ 4.1 EchoMod: Instance Modification
Motivation  When labels are noisy, they do not reflect the true characteristics of the corresponding data instances. This discrepancy hinders a model’s ability to learn meaningful patterns. EchoMod addresses this by transforming data instances to be consistent with their noisy labels. This controlled modification helps the model extract relevant information even when labels contain noise.
@@
Mechanism  EchoMod leverages a pre-trained controllable generative model (e.g., a controllable diffusion-based model) to modify data instances. The primary goal is to enhance the alignment between instances and their potentially noisy labels. This alignment is achieved by carefully guiding the generative model’s process. First, the controllable generative model has undergone prior training on a large dataset. This pre-training has equips the model with a deep understanding of the patterns and structures inherent in the data domain. Second, EchoMod provides both the original instance ($X$) and the noisy label ($\tilde{Y}$) as inputs to the generative model. This dual conditioning shapes the output, encouraging the model to produce a modified instance ($X^{\prime}$) that closely aligns with the noisy label while still preserving essential characteristics of the original data. Striking this balance between label alignment and preventing excessive distortion is crucial.
@@
Effectiveness and Flexibility  Handling label noise in noisy label learning is a well-recognized challenge. Previous works have primarily focused on the utilization and optimization of internal data. Our approach introduces a novel perspective by integrating external knowledge to enhance model robustness. This integration does not compromise fairness, as our flexible framework accommodates various generative models, and can be fine-tuned for specific noisy label problems. By doing so, we ensure that the method does not overly rely on any particular model. This approach is particularly advantageous when dealing with ambiguous data. The inherent ambiguity in the data can lead to low confidence in direct discrimination. Instead, by modifying the input data through controllable generative models, we can better resolve discrepancies between instances and labels while preserving the meaningful characteristics of the original data. This not only improves the model’s discriminative capability but also enhances overall performance and reliability.
@@
Flexible Generalization with Minimal Tuning  In most cases, EchoMod can leverage a pre-trained controllable generative model without fine-tuning during the alignment process. This preserves the model’s ability to understand general data characteristics, promoting EchoAlign’s applicability across various domains. While EchoMod can be effective without fine-tuning, additional performance gains might be realized by tailoring the controllable generative model to highly specialized tasks or data distributions. In such cases, fine-tuning could lead to better alignment between instances and noisy labels, especially in specialized applications such as medical imaging or scientific data.
Visualization and Comparison of Generation Examples  To further illustrate the significance of controllable generative models (CGM) and their advantages over non-controllable generative models (NGM), we provide additional generation examples in Table 1. Specifically, we compare the modifications produced by two representative controllable generative models (ControlNet and UniControl) against two advanced non-controllable generative models (GPT-4 and Gemini).
As depicted, CGMs successfully maintain intrinsic instance characteristics while aligning instances effectively with their noisy labels. In contrast, NGMs struggle to preserve crucial features, often leading to semantic misalignments or unrealistic modifications. For instance, when converting a hoodie to a T-shirt, CGMs effectively adjust clothing style while preserving facial and body features, whereas NGMs drastically distort or remove essential details.
These examples empirically validate our theoretical analysis (§ 3), demonstrating that CGMs substantially enhance instance-label alignment, reduce information distortion, and facilitate stable predictions in noisy label scenarios. Thus, controllable generative models emerge as a more robust and reliable choice for instance modification tasks.
CGM NGM
Noisy Label Original Instance ControlNet UniControl GPT-4 Gemini
@@ 4.2 EchoSelect: Instance Selection
Motivation  While EchoMod improves the alignment between instances and noisy labels, some modified instances may still exhibit inconsistencies. Additionally, the instance modification process can introduce distribution shifts between the modified training data and the true test distribution. EchoSelect safeguards against these issues by identifying and retaining only the most reliable instances after modification. This filtering enhances model robustness, reduces the impact of noisy data, and mitigates distribution shifts introduced by instance modification.
Mechanism  EchoSelect employs a metric to assess the similarity between modified instances and a reference representation of clean data. We use the Cosine similarity between feature vectors extracted using a suitable feature extractor (e.g., the image encoder of CLIP (Radford et al.,, 2021)):
@@
Selection Process  EchoSelect calculates similarity for all modified instances, comparing them to their original counterparts. To mitigate distribution shifts, priority is given to maintaining clean original instances as much as possible. The final training set consists of two parts: (1) Original instances with similarity above a determined threshold are deemed sufficiently aligned with clean data characteristics and retained, and (2) modified instances with similarity below the threshold are included. These instances are likely those where the modification was most beneficial in aligning them with the noisy labels, while also indicating some degree of difference from the original distribution. The threshold $\tau$ balances the inclusion of modified instances with the preservation of the original data distribution, ensuring that only instances aligned with the characteristics of clean data are retained. Our sensitivity analysis (§5.3) confirms the robustness of $\tau$ across various types of noise.
@@ 4.3 EchoAlign: Optimized Combination
[…]
@@
[…]
2: Compute similarity: $S(X^{\prime},X)$ using Equation equation 1
@@
The integration of EchoMod and EchoSelect enables the creation of a refined training dataset that is aligned with noisy labels and filtered for quality. This optimized dataset is better suited for robust learning in the presence of large label noise. Since the refined training dataset can be further used to train a supervised or self-supervised model for LNL, EchoAlign can be combined with advanced LNL methods to further mitigate the impact of label noise. The integration of EchoMod and EchoSelect is encapsulated in Algorithm 1, which details the steps for modifying instances and selecting the optimal subset.
@@ 5 Experiments
To evaluate the robustness and effectiveness of our proposed method, we conducted a comprehensive set of experiments across multiple datasets and baseline comparisons. The detailed implementation settings, including model configurations, hyperparameters, and data preprocessing, are provided in Section 5.1.
@@ 5.1 Experiment Setup
Dataset  Our experiments are conducted on two synthetic datasets: CIFAR-10 and CIFAR-100 (Krizhevsky et al.,, 2009), and a real-world dataset: CIFAR-10N (Wei et al.,, 2022). CIFAR-10 and CIFAR-100 each contain 50,000 training and 10,000 testing images, with a size of $\times32$, covering 10 and 100 classes respectively. CIFAR-10N utilizes the same training images from CIFAR-10 but with labels re-annotated by humans. Although CIFAR-10 is a clean dataset, inherent ambiguity in many images leads to prevalent label noise, as even humans struggle to provide consistent labels, a phenomenon reflected in CIFAR-10N. Following previous research protocols (Bai et al.,, 2021; Xia et al.,, 2019; Xia et al., 2023b, ), we corrupted these synthetic datasets using three types of label noise. Specifically, symmetric noise randomly alters a proportion of labels to different classes to simulate random errors; pair flip noise changes labels to adjacent classes with a certain probability; and instance-dependent noise modifies labels based on image features to related incorrect classes. Due to the inherent ambiguity in CIFAR-10N images, correcting label noise has limited impact on performance, making it a more practical choice over Clothing1M (Xiao et al.,, 2015). A detailed runtime analysis is provided in Section 5.4, demonstrating that the runtime is reasonable across different datasets and can be further optimized using model acceleration techniques. For CIFAR-10N, we use four noisy label sets: ‘Random i=1, 2, 3’, each representing the label provided by one of three independent annotators; and ‘Worst’, which selects the noisiest label when incorrect annotations are present.
@@
Baseline  We compare EchoAlign against various paradigms of baselines for addressing label noise. Under the robust loss function paradigm, we include APL (Ma et al.,, 2020), PCE (Menon et al.,, 2019), AUL (Zhou et al.,, 2023), and CELC (Wei et al.,, 2023); under the loss correction paradigm, we adopt T-Revision (Xia et al.,, 2019) and Identifiability (Liu et al.,, 2023); under the label correction paradigm, we select Joint (Tanaka et al.,, 2018); and under the sample selection paradigm, we employ Co-teaching (Han et al.,, 2018), SIGUA (Han et al.,, 2020), and Co-Dis (Xia et al., 2023a, ). We compare these methods against a simple cross-entropy (CE) loss baseline. Following the fair baseline design proposed by Xia et al., 2023b , we do not compare with methods such as MixUp (Zhang et al.,, 2018), DivideMix (Li et al.,, 2020), and M-correction (Arazo et al.,, 2019), as these involve semi-supervised learning, making such comparisons unfair due to inconsistent settings.
@@
Implementation Details  All experiments were conducted on an NVIDIA V100 GPU using PyTorch. The model architectures and parameter settings were kept consistent with previous studies (Bai et al.,, 2021). The experiments were configured with a learning rate of 0.1, using the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9, and a weight decay set to $1\times 10^{-4}$. We applied 30% and 50% symmetric noise and 45% pair flip noise on the CIFAR-10 and CIFAR-100 datasets to assess model performance. The CIFAR-10 dataset utilized the standard ResNet-18 (He et al.,, 2016) architecture, while CIFAR-100 used ResNet-34. For the CIFAR-10N dataset, the same ResNet-18 model was used. Prior to training, ControlNet was utilized as our reference model in the controllable generation model module. This choice was strategic; ControlNet was the least effective model identified in prior analyses (Chen et al., 2023c, ). Employing this model underscores the robustness of our approach, ensuring that the efficacy of our method is not overly contingent upon the capabilities of any specific generative model. This decision highlights our method’s adaptability and general efficacy across varying scenarios. We employed the Canny edge detector as a simple preprocessor to extract features from the instances, using labels as textual controls with the prompt "a photo of {label}". No additional or negative prompts were used, and the sampling process was limited to 20 steps. All experiments were repeated three times with different random seeds, and results are reported as averages with standard deviations.
Data preprocessing  For all datasets, including CIFAR-10, CIFAR-100, and CIFAR-10N, we adopted a unified data augmentation strategy. Specifically, we first applied 4-pixel padding, followed by random cropping to $32\times 32$ pixels. We then applied random horizontal flipping and normalization.
Hyperparameter settings  For the ControlNet controllable generation model, we used the simplest Canny preprocessor with both the low threshold and high threshold set to 75. The prompt used was “a photo of label” without any additional prompts or negative prompts. The feature maps output by the preprocessor and the generated images both had a medium size of $512\times 512$ pixels. The diffusion process consisted of 20 steps. For the EchoSelect section, the default threshold was set to 0.4 for all cases with 30% noise, and 0.52 for cases with 45% and 50% noise. The hyperparameters for the training are detailed in Table 2.
CIFAR-10/CIFAR-10N CIFAR-100
architecture ResNet-18 ResNet-34
optimizer SGD SGD
loss function CE CE
learning rate(lr) 0.1 0.1
lr decay 100th and 150th 100th and 150th
weight decay $10^{-4}$ $10^{-4}$
momentum 0.9 0.9
batch size 128 128
training samples 45,000 45,000
training epochs 200 200
@@
5.2 Main Results
We evaluated our method on two synthetic datasets (CIFAR-10 and CIFAR-100) and one real-world dataset (CIFAR-10N). For CIFAR-10 and CIFAR-100, 90% of the noisy-labeled data was used for training, 10% for validation, and evaluation was performed on clean test samples. Several baseline results were obtained from previous work (Xia et al., 2023b, ). As shown in Table 3, our method achieved state-of-the-art performance in most scenarios. Under challenging noise conditions (e.g., 50% instance-dependent noise on CIFAR-10 and 45% symmetric noise on CIFAR-100), our method significantly outperformed existing baselines, demonstrating its robustness against various types of noise. This robustness is particularly attributable to EchoMod’s noise-independence, which enables the model to learn consistent features across different noise types and levels. Performance variations were mainly caused by differences in the number of clean samples in the datasets. On the real-world CIFAR-10N dataset, our method also outperformed state-of-the-art methods across all noise settings, exhibiting strong robustness with minimal variations in performance.
@@ 5.3 In-Depth Analyses
Ablation Analysis  To assess the effectiveness of EchoMod and EchoSelect, we conducted ablation studies by systematically disabling these components. Specifically, we evaluated two configurations: “Instance Modification Only” and “EchoSelect Only,” and compared both against the standard Cross-Entropy Loss (CE) as a baseline. These experiments were carried out under several settings with high noise rates, presenting significant challenges for the model. The experimental results in Table 5 revealed that when using only Instance Modification, the model’s accuracy did not exceed the baseline CE, and even decreased. This decline primarily stems from the data distribution shift caused by solely using modified instances, adversely affecting the model’s generalization capability. In contrast, using only EchoSelect improved performance but still fell short of the combined EchoAlign approach. This indicates that although EchoSelect significantly reduces the impact of noise, its effectiveness is limited by the number of available samples.
@@
Sensitivity Analysis  The performance of EchoSelect is influenced by the threshold value $\tau$, which affects the number and quality of samples selected from noisy datasets. According to the assumptions of EchoSelect, the optimal threshold should theoretically be around 0.5. To validate the efficacy of our method, we used the state-of-the-art BLTM (Yang et al.,, 2022) approach as a baseline, with results directly cited from its original publication. As shown in Table 6, EchoSelect was able to select significantly more samples than BLTM under the same accuracy conditions. Particularly, under 30% instance-dependent noise, when the accuracy reached 99%, EchoSelect retained almost twice as many samples as BLTM.